OpenStudy (anonymous):

What does 4cosΘ^2 simplify to?? help

OpenStudy (anonymous):

The question was (cos Θ + cos Θ)2 + (cos Θ + cos Θ)2 and my options are sin2 Θ 4 8cos2 Θ cos2 Θ

OpenStudy (anonymous):

the 2 in the question are exponents

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (\cosθ+\cosθ)^2+(\cosθ+\cosθ)^2 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ (2\cosθ)^2+(2\cosθ)^2 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 4\cos^2θ+ 4\cos^2θ \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 8\cos^2θ \LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

@SolomonZelman ahh i made a dumb mistake thank you sooo much!! can u help me with one more? its on graphing sequences

OpenStudy (anonymous):

wait, for the one you just answered why does it go from 2cos(theta)^2 to 4cos(theta)^2?

OpenStudy (solomonzelman):

graphing sequences? Well... you can write them as coordinates and graph. Geometric sequences are curves (exponential functions), Arithmetic sequences are lines, (Linear functions).

OpenStudy (solomonzelman):

Because you raise all the parenthesis to the second power, and the `2` is being raised to the second power also.

OpenStudy (anonymous):

@SolomonZelman oh okay! i was a bit confused

OpenStudy (solomonzelman):

x coordinate is the number of terms, and the y coordinates are the terms themselves.

OpenStudy (solomonzelman):

Can you find the common ratio ?

OpenStudy (anonymous):

but theres no zeroth term?

OpenStudy (anonymous):

one sec i think i can get it

OpenStudy (anonymous):

2 right? @solomonzelman

OpenStudy (solomonzelman):

yes there is no zeroth term, and you have no intercepts...

OpenStudy (solomonzelman):

the common ratio can not be 2, because the terms get smaller. 10, 5, 1.25 ...

OpenStudy (anonymous):

oh is it 1/2?

OpenStudy (solomonzelman):

I mean 10, 5, __, 1.25

OpenStudy (solomonzelman):

Yes it is 1/2

OpenStudy (solomonzelman):

So, can you find the average rate ?

OpenStudy (anonymous):

but from what? like between what two points

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Like you find the overall rate, between the 4 points, and since you have 3 distances from 1st term to the second from second to the 3rd (and) from 3rd to the 4th you divide the overall distance by 3 to get the average rate.

OpenStudy (anonymous):

ohh okay one sec ill do that

OpenStudy (anonymous):

I got: -5 -2.5 -.125

OpenStudy (anonymous):

ive narrowed down the choices to b and d since 10 is the first term of the sequence

OpenStudy (anonymous):

i just don't know if the average rate of change is negative or not. i think it is because the numbers i just got are all negative

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ d=\sqrt{(10-1.25)^2+(1-4)^2} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ d=\sqrt{(8.75)^2+(-3)^2} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ d=\sqrt{85.5625 } \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ d=9.25 \LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 9.25=9\frac{1}{4} \LARGE\color{white}{ \rm │ }}\)

OpenStudy (anonymous):

is 9.25 the average rate of change?

OpenStudy (solomonzelman):

that is the distance (in a straight line) from x=1 to x=4.

OpenStudy (anonymous):

oh okay. but how do i determine if the average rate of change is 15/2 or -15/2?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

What do you think? is it a negative or a positive rate of change ?

OpenStudy (solomonzelman):

(x → ∞, y→0 )

OpenStudy (anonymous):

negative

OpenStudy (solomonzelman):

Yes

OpenStudy (anonymous):

yay!! thank you so much i appreciate it

OpenStudy (solomonzelman):

you answer, is choice ?

OpenStudy (solomonzelman):

( just to verify )